OpenStudy (anonymous):

I have this homework and I need to find the formulae for S_n in function of "n" The sum is sum (3)/4^(n+1) n=1 to infinity. I have calculated the first four terms of the series. These are the first four: 3/16, 3/64,3/256, 3/1024. According my book I have to do: S1=u1 and S2=S1+u2 S3=S2+u3 S4=S3+u4 As uk=34k+1 Rewriting I got: uk=34k4 or 3∗4−k−1. Am I right?? What I can do more from here?

OpenStudy (anonymous):

I´m searching for this answer S_n=(4^(n)-1)/4^(n+1).

OpenStudy (phi):

like this: \[ \frac{4^n -1}{4^{n+1} } \] ?

OpenStudy (anonymous):

And What more?? How can I get to the answer?? Now I know that S_1=u_1 and S_2=S_1+u_2 S_3=S_2+u3 S_4=S_3+u_4 According to the book u_k=3/4^(k+1) And by book´s example I have to write: S_n=u_1+u_2+u_3+u_4+...u_n-1+u_n. I know that the discretized function that is U_n is (3)/4^(n+1). In this case U_n-1 it will be (-3)/4^(-n-1)?

OpenStudy (anonymous):

And What more?? How can I get to the answer?? Now I know that S_1=u_1 and S_2=S_1+u_2 S_3=S_2+u3 S_4=S_3+u_4 According to the book u_k=3/4^(k+1) And by book´s example I have to write: S_n=u_1+u_2+u_3+u_4+...u_n-1+u_n. I know that the discretized function that is U_n is (3)/4^(n+1). In this case U_n-1 it will be (-3)/4^(-n-1)?